How many molecules of O2 are present in a 3.90L flask at a temperature of 273K and a pressure of 1.00atm?

  • Use 6.022×10^23 mol^ −1 for Avogadro's number.

Select the correct answer below:

A) 1.05×10^23 molecules

B) 1.74×10^23 molecules

C) 3.74×10^24 molecules

D) 9.60×10^22 molecules
Answer #1

Answer : option(A) is the correct option

Given : V = 3.9 L , T = 273 K , P = 1atm

And we know R (gas constant ) = 0.0821 L atm mol​​​​-1 K​​​​​​-1

Using ideal gas equation : PV = n * R * T

Where n = no. of moles of O​​​​​​2 gas

n= P * V / (R* T) = 1 * 3.9 /(0.0821 * 273 ) = 0.174 moles

As 1 mole of O​​​​​​2 contain= 6.022 * 1023 molecules

So, 0.174 mole of O​​​​​​2 contain = 0.174 * 6.022 * 1023

= 1.05 * 1023 molecules

Which is option (A)